# Self interacting scalar field theory in general curved spacetimes at zero and finite temperature revisited

### v1

@@ 7 Conclusion
We have computed in 3, the one loop effective potential at zero temperature, up to the quadratic order of the spacetime curvature for quartic and cubic self interactions. In 4, we extended these results at finite temperature, with high temperature approximation. We discuss spontaneous symmetry breaking, its restoration and phase transition at zero and finite temperatures in 5, 5.1. We next generalise some of these results to two loop in 6. In particular, the contribution from the ${\cal O}(\lambda)$ two loop double bubble diagram under local approximation to the effective potential, at finite temperature has been explicitly presented, 6.1. For the de Sitter spacetime in particular, we have shown that we can have symmetry breaking for a scalar even with a positive rest mass squared and a positive non-minimal coupling, at zero temperature, at both one and two loop level. This cannot be achieved by linear curvature approximation and the result remains valid for a very large range of renormalisation scale, even though $\Lambda$ is taken to be as tiny as its current observed value.
@@
The above results perhaps indicate that we should employ some non-perturbative techniques in order to investigate further the symmetry breaking and phase transition phenomena in curved spacetimes. Note that this not only means resummation of the coupling constant, but also at least some of the curvature terms appearing in the expansion of the propagator. This proposition is also motivated by the one loop self energy computation, 27. Similar analyses would be highly interesting for (non-)Abelian gauge theories. We hope to come back to this issue sometimes soon in a different publication.
@@ Acknowledgement
The work of VN is supported by the junior research fellowship of University Grants Comission, Govt. of India (Ref. no. 211610068936). He also wishes to thank B. L. Hu and I. L. Shapiro for useful communications.

### v2

@@ 7 Conclusion
We have computed in 3, the one loop effective potential at zero temperature, up to the quadratic order of the spacetime curvature for quartic and cubic self interactions. In 4, we extended these results at finite temperature, with high temperature approximation. We discuss spontaneous symmetry breaking, its restoration and phase transition at zero and finite temperatures in 5, 5.1. We next generalise some of these results to two loop in 6. In particular, the contribution from the ${\cal O}(\lambda)$ two loop double bubble diagram under local approximation to the effective potential, at finite temperature has been explicitly presented, 6.1. For the de Sitter spacetime in particular, we have shown that we can have symmetry breaking for a scalar even with a positive rest mass squared and a positive non-minimal coupling, at zero temperature, at both one and two loop level. This result remains valid for a very large range of renormalisation scale, and it cannot be achieved by linear curvature approximation. We believe this result to be interesting in its own right. It is clear that due to symmetry breaking in such potential, the new mass of the field will get contributions from the spacetime curvature, as well as non-minimal coupling. We have argued that such effect can be expected to be significant in an early universe scenario, where the density of the dark energy was supposed to be rather high compared to that of today.
As we also noted earlier, the Schwinger-DeWitt expansion of the Feynman propagator, 2, is essentially meant to capture the curvature effects into the ultraviolet of high energy quantum processes, such as the trace anomaly [6]. In other words, this method may not be suitable to probe, e.g., the deep infrared non-equilibrium effects relevant to the late time, super-Hubble scale during the primordial inflation [88, 89, 90, 91, 92, 93, 94] (also references therein), found by using the Schwinger-Keldysh closed time path formalism. Thus the results presented here can be thought of as sub-Hubble and short time scale, transient or even some initial phenomenon. An essential manifestation of this will be to note that the effective potential we found for the de Sitter will also be valid for the static or any other patch, provided we erect our normal coordinate system in a small neighbourhood located much inside the cosmological event horizon. This is because the expansion coefficients of the Green function in 10 consist only of curvature invariants, which are independent of any coordinate system. Probing local Lorentz invariant physical processes in a curved spacetime background, to the best of our knowledge and understanding, are important in their own right. Also, after the occurrence of the SSB at short spacetime scales, the field acquires a new mass, which can affect its dynamics at large scales at late times. It may be interesting to look into the late time long wavelength or stochastic behaviour of the background scalar field, the equation of motion of which can be found from the effective action corresponding to the effective potential.
In this connection we also note that, in a curved spacetime, the notion of temperature and thermal equilibrium might be ambiguous [97]. This may correspond to, for example, the absence of spatial isotropy, the red or blueshift effects etc. However, suppose we confine ourselves to a small region of the spacetime and look for the leading departure from that of the flat spacetime. In that case, it seems that such issues are taken care of, at least approximately. Certainly by construction, the Schwinger-DeWitt expansion technique seems to be appropriate to address such local thermal issues in a curved background, e.g. [74].
@@
Also, as we noted earlier towards the end of 5, that the above results perhaps indicate that we should employ some non-perturbative techniques in order to investigate further the symmetry breaking and phase transition phenomena in curved spacetimes. This not only means resummation of the coupling constant, but also at least some of the curvature terms appearing in the expansion of the propagator. This proposition is also motivated by the one loop self energy computation, 28. Note once again that such resummed formalism is not expected to capture the infrared physics, as the framework would still be based upon the local Lorentz invariance and four momenta. Instead, this attempts to resum the curvature effects in the local, ultraviolet phenomenon. Finally, we also note that the $\phi^{4}$ theory is a toy model, and in order to see whether there is any phenomenological consequences of the results we obtain here, we must go to scalars coupled to non-Abelian gauge fields. While the present work can be considered as a first step in order to gain some initial insights, we hope to come back to the aforementioned issues sometimes soon in future publications.
@@ Acknowledgement
The work of VN is supported by the junior research fellowship of University Grants Comission, Govt. of India (Ref. no. 211610068936). He also wishes to thank B. L. Hu and I. L. Shapiro for useful communications. The authors would like to acknowldge anonymous referees for various useful comments.